# A variant of the Linnik-Sprindžuk theoremfor simple zeros of Dirichlet $L$-functions

### v1

@@ Abstract.
For a primitive Dirichlet character ${\mathfrak{X}}$, a new hypothesis ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is introduced, which asserts that (1) all simple zeros of $L(s,{\mathfrak{X}})$ in the critical strip are located on the critical line, and (2) these zeros satisfy some specific conditions on their vertical distribution. Hypothesis ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is likely to be true since it is a consequence of the generalized Riemann hypothesis.
@@
Keywords: generalized Riemann hypothesis, Riemann zeta function, simple zeros.
@@ 1. Introduction
holds as $x\to 0^{+}$ for any coprime integers $h,k$ with $20<|h|\leqslant k/$ 2, where the sum runs over the imaginary parts $\gamma$ of the nontrivial zeros of $\zeta(s)$. This is known as the Linnik–Sprindžuk theorem. Some similar results have been obtained by Fujii [6, 7, 8], Suzuki [15], Kaczorowski and Perelli [10], and the author [1, 2, 3].
In the present paper, we establish a variant of the Linnik–Sprindžuk theorem focused on the simple zeros of Dirichlet $L$-functions. Assuming the generalized Lindelöf hypothesis, we show that the horizontal and vertical distribution of the simple zeros of any single $L$-function $L(s,\chi)$ can strongly influence the horizontal and vertical distribution of the simple zeros of any other Dirichlet $L$-function. In particular, we state a new criterion for the nonexistence of Siegel zeros.
Our main results are formulated in §2 after the necessary notation has been introduced. To prove the main theorem, we study the interplay and consequences of the following hypotheses on $L$-functions attached to Dirichlet characters $\chi$.
@@
Hypothesis ${\tt RH}_{\tt sim}[\boldsymbol{\star}]$: The hypothesis ${\tt RH}_{\tt sim}[\chi]$ holds for every character $\chi$.
These two hypotheses lie at the heart of our work. ${\tt RH}_{\tt sim}[\chi]$ is a weak form of ${\tt RH}[\chi]$ which asserts that the simple zeros of $L(s,\chi)$ in the critical strip all lie on the critical line; nothing is assumed about trivial zeros of $L(s,\chi)$ (which are all simple) or zeros of multiplicity two or more in the critical strip (that is, non-simple zeros).
@@
The most important hypothesis from the perspective of the present paper, namely ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$, can only be formulated after the necessary notation has been introduced; see §2.
@@ 2. Notation and statement of results
For an arbitrary character $\chi$, we make extensive use of the function $D(s,\chi)$, which is defined in the half-plane $\{\sigma>1\}$ by
@@
Note that $0\leqslant\ell(n)\leqslant(\log n)^{2}$. The function $D(s,\chi)$ is meromorphic with a simple pole of residue $L^{\prime}(\rho,\chi)$ at every simple zero $\rho$ of $L(s,\chi)$, and a pole of order three at $s=1$ when $\chi$ is principal. On the other hand, $D(s,\chi)$ is analytic in a neighborhood of any non-simple zero $\rho$ of $L(s,\chi)$.
@@
The next result is a variant of Conrey, Ghosh, and Gonek [4, Lemma 1]. The proof relies on Gonek [9, Lemma 2]; it is based on the stationary phase method.
@@
$g(s)\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=\prod_{p\,\mid\,q}(1-p^{-s}),$
@@
where the sum runs over simple zeros of $L(s,{\mathfrak{X}})$ in the critical strip ((((counted with multiplicity)))),
@@
Remarks.
Hypothesis ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is true for all primitive characters ${\mathfrak{X}}$ under the generalized Riemann hypothesis. It is important to observe that ${\tt RH}^{\dagger}[\chi]$ has been formulated entirely in terms of the simple zeros of a single $L$-function $L(s,{\mathfrak{X}})$; its validity depends only on the horizontal and vertical distribution of the simple zeros. If ${\mathfrak{X}}$ and ${\mathfrak{X}}_{\circ}$ are different primitive characters, then there is no reason, a priori, that the hypotheses ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ and ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}_{\circ}]$ should be related to one another. Nevertheless, we show that these hypotheses are equivalent if one assumes ${\tt LH}[\boldsymbol{\star}]$.
Our main results are the following.
@@
Assume ${\tt LH}[\boldsymbol{\star}]$. For each primitive character ${\mathfrak{X}}$, the hypotheses ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ and ${\tt RH}_{\tt sim}[\boldsymbol{\star}]$ are equivalent. Thus, If ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ holds for some primitive character ${\mathfrak{X}}$, then it holds for every primitive character ${\mathfrak{X}}$.
@@
Assume ${\tt LH}[\boldsymbol{\star}]$. If there is a primitive character ${\mathfrak{X}}$ such that ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is true, then Siegel zeros do not exist.
Indeed, under the conditions of the corollary, Theorem 2.3 shows that ${\tt RH}_{\tt sim}[\boldsymbol{\star}]$ is true. Thus, for every character $\chi$, all simple zeros of $L(s,\chi)$ in the critical strip must lie on the critical line.
@@ 4. Criteria for ${\tt RH}_{\tt sim}[\chi]$
$(i)\Longrightarrow(ii)$. Let $T\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=\sqrt{X}$, $\sigma_{0}\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=1+\frac{1}{\log X}$, and $\sigma_{1}\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=\frac{1}{2}+\frac{1}{\log X}$. Let $\mathcal{C}$ be the rectangular contour in $\mathbb{C}$ consisting of the four directed line segments:

### v2

@@ Abstract.
For a primitive Dirichlet character ${\mathfrak{X}}$, a new hypothesis ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is introduced, which asserts that (1) all simple zeros of $L(s,{\mathfrak{X}})$ in the critical strip are located on the critical line, and (2) these zeros satisfy some specific conditions on their vertical distribution. We show that ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ (for any ${\mathfrak{X}}$) is a consequence of the generalized Riemann hypothesis.
@@
Keywords: generalized Riemann hypothesis, Riemann zeta function, Lindelöf hypothesis, simple zeros.
@@ 1. Introduction
holds as $x\to 0^{+}$ for any coprime integers $h,k$ with $20<|h|\leqslant k/$ 2, where the sum runs over the ordinates $\gamma$ of the nontrivial zeros of $\zeta(s)$. This is known as the Linnik–Sprindžuk theorem. Some similar results have been obtained by Fujii [6, 7, 8], Suzuki [15], Kaczorowski and Perelli [10], and the author [1, 2, 3].
In the present paper, we establish a variant of the Linnik–Sprindžuk theorem focused on the simple zeros of Dirichlet $L$-functions. Assuming the generalized Lindelöf hypothesis, we show that the horizontal and vertical distribution of the simple zeros of any single $L$-function $L(s,\chi)$ can strongly influence the horizontal and vertical distribution of the simple zeros of any other Dirichlet $L$-function. In particular, we give a new criterion for the nonexistence of Siegel zeros.
Our results are formulated in §2 after the necessary notation has been introduced. To prove the main theorem, we study the interplay and consequences of various hypotheses on the $L$-functions attached to Dirichlet characters $\chi$.
@@
Hypothesis ${\tt RH}_{\tt sim}[\boldsymbol{\star}]$: Hypothesis ${\tt RH}_{\tt sim}[\chi]$ is true for every character $\chi$.
The latter two hypotheses lie at the heart of our work. ${\tt RH}_{\tt sim}[\chi]$ is a weak form of ${\tt RH}[\chi]$ which asserts that the simple zeros of $L(s,\chi)$ in the critical strip all lie on the critical line; nothing is assumed about the trivial zeros of $L(s,\chi)$ (which are all simple) or any zeros of multiplicity two or more in the critical strip (that is, non-simple zeros).
@@
The most important hypothesis from the perspective of the present paper, namely ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$, can only be formulated after some notation has been introduced; see §2 below.
@@ 2. Notation and statement of results
For an arbitrary character $\chi$, we make extensive use of the function $D(s,\chi)$ defined in the half-plane $\{\sigma>1\}$ by
@@
Note that $0\leqslant\ell(n)\leqslant(\log n)^{2}$. The function $D(s,\chi)$ is meromorphic with a simple pole of residue $L^{\prime}(\rho,\chi)$ at every simple zero $\rho$ of $L(s,\chi)$, and a pole of order three at $s=1$ in the case that $\chi$ is principal. On the other hand, $D(s,\chi)$ is analytic in a neighborhood of any non-simple zero $\rho$ of $L(s,\chi)$.
@@
The next result is a variant of Conrey, Ghosh, and Gonek [4, Lemma 1]. The proof relies on Gonek [9, Lemma 2] and is based on the stationary phase method.
@@
$g_{q}(s)\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=\prod_{p\,\mid\,q}(1-p^{-s}),$
@@
where the sum runs over simple zeros of $L(s,{\mathfrak{X}})$ in the critical strip,
@@
Observe that ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is formulated entirely in terms of the simple zeros of a single $L$-function $L(s,{\mathfrak{X}})$; the hypothesis asserts nothing about any zeros of higher order that might exist. For characters ${\mathfrak{X}}_{1}\neq{\mathfrak{X}}_{2}$, there is no reason a priori to expect that the hypotheses ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}_{1}]$ and ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}_{2}]$ are related. However, our principal theorem reveals a close connection.
@@
Assume ${\tt LH}[\boldsymbol{\star}]$. For each primitive character ${\mathfrak{X}}$, the hypotheses ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ and ${\tt RH}_{\tt sim}[\boldsymbol{\star}]$ are equivalent.
Thus, assuming ${\tt LH}[\boldsymbol{\star}]$, Theorem 2.3 implies that ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}_{1}]\Longleftrightarrow{\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}_{2}]$ for all primitive characters ${\mathfrak{X}}_{1}$ and ${\mathfrak{X}}_{2}$.
@@
Under GRH, ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ holds for every primitive character ${\mathfrak{X}}$.
This is clear since GRH implies both ${\tt LH}[\boldsymbol{\star}]$ and ${\tt RH}_{\tt sim}[\boldsymbol{\star}]$.
Corollary 2.5.
Assume ${\tt LH}[\boldsymbol{\star}]$, and suppose ${\tt RH}_{\tt sim}^{\dagger}[{\mathfrak{X}}]$ is true for some primitive character ${\mathfrak{X}}$. Then the simple zeros of any Dirichlet $L$-function $L(s,\chi)$ lie on the critical line. In particular, Siegel zeros do not exist.
@@ 4. Criteria for ${\tt RH}_{\tt sim}[\chi]$
$(i)\Rightarrow(ii)$. Let $T\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=\sqrt{X}$, $\sigma_{0}\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=1+\frac{1}{\log X}$, and $\sigma_{1}\mathrel{\vbox{\hbox{\scriptsize.} \hbox{\scriptsize.}}}=\frac{1}{2}+\frac{1}{\log X}$. Let $\mathcal{C}$ be the rectangular contour in $\mathbb{C}$ consisting of the four directed line segments: